# Learning stochastic dynamics from snapshots through regularized unbalanced optimal transport

### v3

@@ 4 Regularized Unbalanced Optimal Transport
where $\Psi:\mathbb{R}\rightarrow[0,+\infty]$ corresponds to the growth penalty function, the infimum is taken over all pairs $\left(p,\boldsymbol{b}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
@@
$\inf_{\left(p,\boldsymbol{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g+\alpha\Psi\left(g\right)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (7)
@@
Here $\boldsymbol{v}(\boldsymbol{x},t)$ represents a new vector field. The proof is left to Section D.2. In (Baradat & Lavenant, 2021), they proposed another Fisher regularization form of RUOT, expressed in the formula Eq. 15. In their formulation, computing the cross term $\langle\nabla_{\boldsymbol{x}}\log p,\sigma^{2}(t)\boldsymbol{v}\rangle$ required calculating the derivative of the function $\log p$ with respect to $\boldsymbol{x}$ and performing vector multiplications with $\boldsymbol{v}$. The formulation here is equivalent to theirs but more computationally tractable, as we avoid differentiation and vector multiplication in our cross-term $\sigma^{2}(t)(1+\log p)g/2$.
@@
When $g=0$ and $\Psi(0)=0$, then Eq. (7) is the same as the dynamic entropy-regularized optimal transport form as discussed in (Bunne et al., 2023a; Li et al., 2020; Gentil et al., 2017; Pooladian & Niles-Weed, 2024; Neklyudov et al., 2024; Léger & Li, 2021).
@@ 5.1 Energy Loss
$\mathcal{L}_{\text{Energy}}=$ $\mathbb{E}_{\boldsymbol{x_{0}}\sim p_{0}}\int_{0}^{T}\left[\frac{1}{2}\left\|\boldsymbol{v}_{\theta}(\boldsymbol{x}(t),t)\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\boldsymbol{x}}s_{\theta}\right\|_{2}^{2}-\left(\frac{\sigma^{2}(t)}{2}+s_{\theta}\right)g_{\theta}+\alpha\Psi\left(g_{\theta}\right)\right]w_{\theta}(t)\mathrm{d}t,$ (10)
@@ 5.3 Fokker-Planck Constraint and Two Stage Training
9:$\mathcal{L}_{\text{Energy}}\leftarrow\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{t}}\int_{t}^{t+1}\left[\frac{1}{2}\left\|\boldsymbol{v}_{\theta}\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\boldsymbol{x}}s_{\theta}\right\|_{2}^{2}-\left(\frac{\sigma^{2}}{2}+s_{\theta}\right)g_{\theta}+\alpha\Psi\left(g_{\theta}\right)\right]{w}(z)\mathrm{d}z$
@@ 6 Numerical Results
This section evaluates the DeepRUOT solver’s ability to recover growth and transitions accurately and produce realistic stochastic dynamics for constructing a Waddington’s developmental landscape. We take $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$ in the following computations.
@@ D.2 Proof of Fisher Regularization of RUOT
$H\left(p(\boldsymbol{x},1)\right)-H\left(p(\boldsymbol{x},0\right)$ $=\int_{0}^{1}\partial_{t}H\left({p}(\boldsymbol{x},t)\right)\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\left(1+\log{p}(\boldsymbol{x},t)\right)\partial_{t}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\left(1+\log{p}(\boldsymbol{x},t)\right)\cdot\left(-\nabla_{\boldsymbol{x}}\cdot\left({p}(\boldsymbol{x},t)\boldsymbol{v}(\boldsymbol{x},t)\right)+g{p}(\boldsymbol{x},t)\right)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}{p}(\boldsymbol{x},t)\left\langle\nabla_{\boldsymbol{x}}\log{p}(\boldsymbol{x},t),\boldsymbol{v}(\boldsymbol{x},t)\right\rangle+\left(1+\log{p}(\boldsymbol{x},t)\right)gp\mathrm{d}\boldsymbol{x}\mathrm{d}t.$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\left\langle\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p,\boldsymbol{v}(\boldsymbol{x},t)\right\rangle\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$ $=\sigma^{2}(t)\left(H\left(p(\boldsymbol{x},1)\right)-H\left(p(\boldsymbol{x},0)\right)\right)$
$-\int_{0}^{1}\int_{\mathbb{R}^{d}}\sigma^{2}(t)\left(1+\log{p}(\boldsymbol{x},t)\right)gp\mathrm{d}\boldsymbol{x}\mathrm{d}t.$
@@ D.3 RUOT with Nonlinear Fokker-Planck Equation Constraints
When $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$, the regularized unbalanced optimal transport problem (7) is equivalent to
$\inf_{\left(p,\boldsymbol{v},\widetilde{g}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}+\alpha|\widetilde{g}(\boldsymbol{x},t)|^{2}-\frac{\sigma^{4}(t)}{16\alpha}(1+\log p)^{2}\right]p\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (16)
@@
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+\widetilde{g}(\boldsymbol{x},t)p+\frac{\sigma^{2}(t)}{4\alpha}(1+\log p)p.$ (17)
@@
$\widetilde{g}=g-\frac{\sigma^{2}(t)}{4\alpha}(1+\log p).$ (18)
@@
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}(t)\boldsymbol{I}p\right)+gp-\frac{\sigma^{2}(t)}{4\alpha}(1+\log p)p.$
@@
$\mathcal{C}\left(\nu_{0},\nu_{1}\right):=\left\{p,\boldsymbol{b},g\left|\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}(t)\boldsymbol{I}p\right)+gp-\frac{\sigma^{2}(t)}{4\alpha}(1+\log p)p\right|\right\},$
@@
Note that when $\sigma(t)$ is constant Definition D.1 is consistent with the form provided in (Buze & Duong, 2023), and it has been conjectured that it is the higher order approximation scheme to the unbalanced optimal transport problem as $\sigma\rightarrow 0$.

### v4

@@ 4 Regularized Unbalanced Optimal Transport
where $\Psi:\mathbb{R}\rightarrow[0,+\infty]$ corresponds to the growth penalty function, the infimum is taken over all pairs $\left(p,\boldsymbol{b},g\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
@@
$\inf_{\left(p,\boldsymbol{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g-\frac{1}{2}\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}\log p+\alpha\Psi\left(g\right)\right]p\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (7)
@@
Here $\boldsymbol{v}(\boldsymbol{x},t)$ represents a new vector field. The proof is left to Section D.2. In (Baradat & Lavenant, 2021), they proposed another Fisher regularization form of RUOT, expressed in the formula Eq. 15. In their formulation, computing the cross term $\langle\nabla_{\boldsymbol{x}}\log p,\sigma^{2}(t)\boldsymbol{v}\rangle$ required calculating the derivative of the function $\log p$ with respect to $\boldsymbol{x}$ and performing vector multiplications with $\boldsymbol{v}$. The formulation here is equivalent to theirs but more computationally tractable, as we avoid differentiation and vector multiplication in our cross-term.
@@
When $g=0$, $\Psi(0)=0$ and $\sigma(t)$ is constant, then Eq. (7) is the same as the dynamic entropy-regularized optimal transport form as discussed in (Bunne et al., 2023a; Li et al., 2020; Gentil et al., 2017; Pooladian & Niles-Weed, 2024; Neklyudov et al., 2024; Léger & Li, 2021).
@@ 5.1 Energy Loss
$\mathcal{L}_{\text{Energy}}=$ $\mathbb{E}_{\boldsymbol{x_{0}}\sim p_{0}}\int_{0}^{T}\left[\frac{1}{2}\left\|\boldsymbol{v}_{\theta}\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\boldsymbol{x}}s_{\theta}\right\|_{2}^{2}-\left(\frac{\sigma^{2}(t)}{2}+s_{\theta}\right)g_{\theta}-\frac{(\sigma^{2}(t))^{\prime}}{\sigma^{2}(t)}s_{\theta}+\alpha\Psi\left(g_{\theta}\right)\right]w_{\theta}(t)\mathrm{d}t,$ (10)
@@ 5.3 Fokker-Planck Constraint and Two Stage Training
9:$\mathcal{L}_{\text{Energy}}\leftarrow\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{t}}\int_{t}^{t+1}\left[\frac{1}{2}\left\|\boldsymbol{v}_{\theta}\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\boldsymbol{x}}s_{\theta}\right\|_{2}^{2}-\left(\frac{\sigma^{2}}{2}+s_{\theta}\right)g_{\theta}-\frac{(\sigma^{2}(t))^{\prime}}{\sigma^{2}(t)}s_{\theta}+\alpha\Psi\left(g_{\theta}\right)\right]{w}(z)\mathrm{d}z$
@@ 6 Numerical Results
This section evaluates the DeepRUOT solver’s ability to recover growth and transitions accurately and produce realistic stochastic dynamics for constructing a Waddington’s developmental landscape. We take $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$ and $\sigma(t)$ is constant in the following computations.
@@ D.2 Proof of Fisher Regularization of RUOT
$\sigma^{2}(1)H\left(p(\boldsymbol{x},1)\right)-\sigma^{2}(0)H\left(p(\boldsymbol{x},0\right)=\int_{0}^{1}\partial_{t}\left(\sigma^{2}(t)H\left({p}(\boldsymbol{x},t)\right)\right)\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}{p}(\boldsymbol{x},t)\log{p}(\boldsymbol{x},t)+\sigma^{2}(t)\left(1+\log{p}(\boldsymbol{x},t)\right)\partial_{t}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\sigma^{2}(t)\left(1+\log{p}(\boldsymbol{x},t)\right)\cdot\left(-\nabla_{\boldsymbol{x}}\cdot\left({p}(\boldsymbol{x},t)\boldsymbol{v}(\boldsymbol{x},t)\right)+g{p}(\boldsymbol{x},t)\right)+\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}{p}(\boldsymbol{x},t)\log{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\sigma^{2}(t){p}(\boldsymbol{x},t)\left\langle\nabla_{\boldsymbol{x}}\log{p}(\boldsymbol{x},t),\boldsymbol{v}(\boldsymbol{x},t)\right\rangle+\sigma^{2}(t)\left(1+\log{p}(\boldsymbol{x},t)\right)gp+\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}{p}(\boldsymbol{x},t)\log{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t.$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\left\langle\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p,\boldsymbol{v}(\boldsymbol{x},t)\right\rangle\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$ $=\left(\sigma^{2}(1)H\left(p(\boldsymbol{x},1)\right)-\sigma^{2}(0)H\left(p(\boldsymbol{x},0)\right)\right)$
$-\int_{0}^{1}\int_{\mathbb{R}^{d}}\sigma^{2}(t)\left(1+\log{p}(\boldsymbol{x},t)\right)gp\mathrm{d}\boldsymbol{x}\mathrm{d}t$
$-\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}{p}(\boldsymbol{x},t)\log{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t.$
@@ D.3 RUOT with Nonlinear Fokker-Planck Equation Constraints
When $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$ and $\sigma(t)=\sigma$ is constant, the regularized unbalanced optimal transport problem (7) is equivalent to
$\inf_{\left(p,\boldsymbol{v},\widetilde{g}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}+\alpha|\widetilde{g}(\boldsymbol{x},t)|^{2}-\frac{\sigma^{4}}{16\alpha}(1+\log p)^{2}\right]p\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (16)
@@
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+\widetilde{g}(\boldsymbol{x},t)p+\frac{\sigma^{2}}{4\alpha}(1+\log p)p.$ (17)
@@
$\widetilde{g}=g-\frac{\sigma^{2}}{4\alpha}(1+\log p).$ (18)
@@
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}\boldsymbol{I}p\right)+gp-\frac{\sigma^{2}}{4\alpha}(1+\log p)p.$
@@
$\mathcal{C}\left(\nu_{0},\nu_{1}\right):=\left\{p,\boldsymbol{b},g\left|\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}\boldsymbol{I}p\right)+gp-\frac{\sigma^{2}}{4\alpha}(1+\log p)p\right|\right\},$
@@
Note that Definition D.1 is consistent with the form provided in (Buze & Duong, 2023), and it has been conjectured that it is the higher order approximation scheme to the unbalanced optimal transport problem as $\sigma\rightarrow 0$.